OpenStudy (yuna1234):

WILL MEDAL! A cylinder has a radius of 10 cm and a height of 8 cm. What is the volume of the cylinder?

OpenStudy (kinged):

The volume of a cylinder is the area of its base (a circle) multiplied by its height.

OpenStudy (yuna1234):

the thing is i am not good with finding volumes with cylinders!

OpenStudy (kinged):

The area of a circle is\[\pi r ^{2}\]

OpenStudy (yuna1234):

yeah

OpenStudy (kinged):

Putting all the parts together, the volume of a cylinder is \[\pi r ^{2} \times h\]

OpenStudy (yuna1234):

ok so do 3.14x10x10x8?

OpenStudy (kinged):

Yep.

OpenStudy (yuna1234):

i actually did do that and the answer that we both got is different than the answer choices...

OpenStudy (kinged):

What are the choices?

OpenStudy (kinged):

Just don't multiply by pi, What's 10*10*8 The answer is in terms of pi. SO you put it as 10pi for example not 31.4...

OpenStudy (yuna1234):

oh!

OpenStudy (yuna1234):

so D?

OpenStudy (kinged):

Yep.

OpenStudy (yuna1234):

Thanks!

OpenStudy (kinged):

You are welcome.